A wave travelling along a string is describe by y(x,t) = 0.35sin(2.75t – 72.5x).


Determine

i. The velocity with which the wave moves along the string


(2 marks)

ii. The displacement of a particle located at x = 21.5 cm and t = 15.5 sec.

(3 marks)


1
Expert's answer
2022-03-10T18:00:43-0500

The general equation of a wave is given by

"y(x,t) = A\sin(\omega t – k x)"

In our case


"y(x,t) = 0.35\sin(2.75t – 72.5x)"

i. The velocity with which the wave moves along the string

"v=\frac{\omega}{k}=\frac{2.75}{72.5}=0.038\:\rm m/s"

ii. The displacement of a particle located at x = 21.5 cm and t = 15.5 sec

"y(0.215,15.5) = 0.35\sin(2.75*15.5 – 72.5*0.215)\\
=0.33\:\rm m"